OpenStudy (hba):

(1+2omega+3omega^2)(3+2omega+omega^2)

OpenStudy (hba):

I have conclude it till 3+15omega+10omega^2+8omega^3

OpenStudy (hba):

Or 11+15omega+10omega^2

hartnn (hartnn):

write 1+2O+3O^2 = 1+O+O^2 + O+2O^2 that would simplify it

hartnn (hartnn):

as 1+O+O^2 =0

OpenStudy (anonymous):

1 + w + w^2 can be equated to zero hence ur left with (w + 2w^2) (2+w) =2w + w^2 +4w^2 +2w^3 now simplify similarly

hartnn (hartnn):

yeah, just like that ^

OpenStudy (anonymous):

w^3 = 1

OpenStudy (hba):

So (1+w+w^2+w+2w^2)(3+2w+w^2)

hartnn (hartnn):

u can simplify it more 1+2O+3O^2 = 1+O+O^2 + O+2O^2 = O+2O^2 = O^2 + (-O)

OpenStudy (hba):

And we know that 1+w+w^2=0 so, (w+2w^2) (3+2w+w^2 ) Right @hartnn

hartnn (hartnn):

yeah do it same for other bracket

OpenStudy (hba):

More over we know that w+w^2 =-1 so, (w+2w^2)(3+w-1) (w+2w^2)(2+w)

OpenStudy (hba):

@hartnn

hartnn (hartnn):

yeah, go ahead

OpenStudy (hba):

2w+w^2+4w^2+2w^3 2w+5w^2+2w^3 2w+2w^3+5w^2 2(w+1)+5w^2 2(-w^2)+5w^2 Therfore the final expression we get is, 3w^2-----Answer :) @hartnn

hartnn (hartnn):

yeah

OpenStudy (hba):

Thanks Alot :)